# Cross-Sectional Dynamics Under Network Structure: Theory and Macroeconomic Applications

### v6

@@ Conclusion
In this paper, I develop the Network-VAR (NVAR); an econometric framework that rationalizes the dynamics of a cross-sectional variable as the dynamic innovation transmission along bilateral links among cross-sectional units. After characterizing dynamics in relation to network connections and inference of both the timing of network effects and the network itself, I consider two applications. First, consistent with an RBC economy with lagged input-output conversion, I use the NVAR to characterize and estimate the latter’s contribution to business cycles, i.e. to persistence in aggregate output growth. Second, I use the NVAR as a dimensionality-reduction technique for forecasting cross-country macroeconomic aggregates.
Rather than assuming time-invariant links, an important methodological step forward would be to develop a tractable framework that links dynamic network effects to dynamic network formation, allowing units to adjust their their connections in reaction to innovations transmitted through the network. This would permit the study of economic growth through network formation in a dynamic context and of the dynamic consequences of network destruction, such as witnessed through geoeconomic fragmentation in international trade. Methodologically interesting in this context is the study of cointegration relationships as a function of network connectedness.
Furthermore, the NVAR could be augmented to accommodate heterogeneous propagation patterns across units or over time. Relatedly, the application of the NVAR for forecasting time series in very high dimensions would benefit from refinements of the crude shrinkage priors used in this paper.
@@ A.4 NVAR($p,q$)
In the previous example, the observed monthly series depends on its value in the past $p$ quarters, i.e. on its value three months ago, six months ago, etc., up to $3p$ months ago.585858 Note that this assumes that transmission happens instantaneously at the end of each (network interaction-) period. Alternative definitions of a smoother transmission inevitably lead towards abandoning the paradigm of discrete time.
Consider next the general case: $q\in\mathbb{Q}_{++}$. We can write $q=q_{1}q_{2}$ with $q_{1}^{-1}\in\mathbb{N}$ and $q_{2}\in\mathbb{N}$.595959 Note that $q_{2}$ is the least common multiple of $q$ and $1$, whereas $q_{1}$ is their greatest common denominator. For example, under monthly observations, $q=4/3$ implies that network interactions occur every three weeks, $q=4/5$ that they occur every five weeks, and $q=30/4$ that they occur every four days. We can model this case as a combination of the $q\in\mathbb{N}$ and $q^{-1}\in\mathbb{N}$ cases: we observe every $q_{2}\in\mathbb{N}$ periods a snapshot of an NVAR($p^{*},1)$ process with $p^{*}=p/q_{1}\in\mathbb{N}$ and with parameters restricted as in the preceding paragraph. For example, under monthly observations and $q=4/3$, this amounts to observing every fourth period a snapshot of a weekly process that depends on its value three weeks ago, six weeks ago, etc.
@@
$\underset{q,G\rightarrow\infty}{\lim}q\frac{\partial y_{t}}{\partial\tilde{\varepsilon}_{tq-G}}=\underset{G\rightarrow\infty}{\lim}\sum_{h=0}^{G}\frac{\partial\tilde{y}_{\tau+G}}{\partial\tilde{\varepsilon}_{\tau+h}}\;.$
@@
which leads to $\text{Cov}(u_{it},u_{jt})=\mathbf{1}\left\{{i=j}\right\}\frac{1}{2}\sigma_{i}^{2}+\frac{1}{4}\alpha_{1}(a_{ij}\sigma_{j}^{2}+a_{ji}\sigma_{i}^{2})+\frac{1}{4}\alpha_{1}^{2}\sum_{k=1}^{n}a_{ik}a_{jk}\sigma^{2}_{k}$. In the former case, $\text{Cov}(u_{it},u_{jt})$ is determined by common exposure to third units, while in the latter case bilateral exposure matters as well. In either case, exposure means first-order connections.616161 To understand the different results for stock and flow variables, assume for simplicity that there are no self-links: $a_{ii}=0\;\forall\;i$. For a stock variable, $y_{it}=\tilde{y}_{i,\tau}$, which implies that $u_{it}$ only depends on $\tilde{u}_{i,\tau}$, not on $\tilde{u}_{i,\tau-1}$: we have $u_{it}=\tilde{u}_{i,\tau}+\alpha_{1}\sum_{j=1}^{n}a_{ij}\tilde{u}_{j,\tau-1}$. As a result, the only (possibly) common terms in $u_{it}$ and $u_{jt}$ are the one-period lagged high-frequency innovations of third units, $\tilde{u}_{k,\tau-1}$, and any comovement between $u_{it}$ and $u_{jt}$ is due to common exposure to these units: $\text{Cov}\left({u_{it},u_{jt}}\right)\neq 0$ iff $\exists\;k$ s.t. $a_{ik},a_{jk}\neq 0$. In contrast, for a flow variable, $y_{it}=\tilde{y}_{i,\tau}+\tilde{y}_{i,\tau-1}$, which implies that $u_{it}$ also depends on $\tilde{u}_{i,\tau-1}$: we have $u_{it}=\tilde{u}_{i,\tau}+\tilde{u}_{i,\tau-1}+\alpha_{1}\sum_{j=1}^{n}a_{ij}\tilde{u}_{j,\tau-1}$. Besides common exposure to third units, comovement is also due to bilateral exposure: $a_{ij}\neq 0\;|\;a_{ji}\neq 0\;\Rightarrow\;\text{Cov}\left({u_{it},u_{jt}}\right)\neq 0$. Under $q=3$, exposure is determined by first- and second-order connections: e.g. for a stock variable, we have $u_{t}=\tilde{u}_{\tau}+\alpha_{1}A\tilde{u}_{\tau-1}+(\alpha_{2}A+\alpha_{1}^{2}A^{2})\tilde{u}_{\tau-2}$, with
@@ Appendix B NVAR: Inference
Relative to the notation used for the Kalman filter, this is with a slight abuse of notation, as $s_{\tau|\tau+1}\neq\mathbb{E}\left[{s_{\tau}|X_{1:\tau},s_{\tau+1}}\right]$ but $s_{\tau|\tau+1}=\mathbb{E}\left[{s_{\tau}|X_{1:\tau},\left[{s_{\tau+1}}\right]_{1}}\right]$, and similarly $P_{\tau|\tau+1}=\mathbb{V}\left[{s_{\tau}|X_{1:\tau},\left[{s_{\tau+1}}\right]_{1}}\right]$. See Kim and Nelson (1999, p. 194) for the adjustments of the CKSS required when $s_{t}$ is a companion form-VAR(1).
@@ C.1 Theory
$\hskip 130.08731pt\text{s.t.}\quad\quad\check{x}_{t,t-2}^{ij}$ $=x^{ij}_{t,t-2}\;,$
@@ C.3 Estimation
The Uniform prior for $\alpha\in[0,1]^{p-1}\cap\{\alpha:\sum_{l=1}^{p-1}\alpha_{l}\leq 1\}$ can be broken up as follows:
@@
For the parameters $(\sigma^{2}_{a},\sigma^{2\prime})^{\prime}\in\mathbb{R}_{++}^{n+1}$, an upper bound has to be chosen so as to specify a proper Uniform prior distribution, since draws from it are needed to initialize the SMC sampler. A high upper bound is desirable so as to avoid domain restrictions in areas with non-trivial likelihood values. However, for efficient sampling, lower values are preferred. I choose a low upper bound to initialize the sampler, but allow the algorithm to disrespect it in search for parameter-values associated with high likelihood values and therefore high posterior mass. Ex-post, I redefine the upper bound as the largest posterior draw among $(\sigma^{2}_{a},\sigma^{2\prime})^{\prime}$ across all models. The resulting MDDs could be re-scaled to take this into account, but this is not necessary, as all models’ MDD is “off” by the same proportionality constant, which means that the ranking is unaffected.676767 Let $s=(\sigma^{2}_{a},\sigma^{2\prime})^{\prime}\in[0,\bar{s}]^{n+1}$. We get the MDD $p(Y)=\int p(Y|\theta_{-s},s)p(\theta_{-s})p(s)d(\theta_{-s},s)=\bar{s}^{-(n+1)}\int p(Y|\theta_{-s},s)p(\theta_{-s})d(\theta_{-s},s)\;.$ If the upper bound is ex-post rescaled to $s^{*}$, the MDD needs to be corrected by multiplying by $(\bar{s}/s^{*})^{n+1}$. The analogous holds with heterogeneous upper bounds. I set $\bar{\sigma}^{2}_{i}=0.5\mathbb{V}[\Delta\tilde{y}_{it}]$ and $\bar{\sigma}^{2}_{a}=\max_{i}\bar{\sigma}^{2}_{i}$.

### v7

@@ Conclusion
In this paper, I develop the Network-VAR (NVAR) – an econometric framework that rationalizes the dynamics of a cross-sectional variable by the dynamic innovation transmission along bilateral links among cross-sectional units – and I consider two applications. First, I estimate the contribution of lagged input-output conversion to business cycles through the lens of an RBC economy, thereby providing a structural underpinning of the NVAR. Second, I use the NVAR as a dimensionality-reduction technique for forecasting cross-country macroeconomic aggregates.
More work is needed to explore the extent to which the NVAR is useful to address sparse factor-issues and improve upon the forecasting performance of alternative dimensionality-reduction techniques. Its application for forecasting very high-dimensional processes would benefit from refinements of the crude shrinkage priors used in this paper.
Furthermore, rather than assuming time-invariant links, an important methodological step forward would be to jointly study dynamic network effects and -formation.3535todo: 35cite Bykhovskaya, P007 and also KuersteinerPrucha20203636todo: 36mention macro studies on growth via network changes? Rather than assuming stationarity, an interesting avenue for further research would be to study networked cointegration. Finally, the NVAR could be augmented to accommodate heterogeneous propagation patterns across units or over time, which in the context of a production economy amounts to endogenizing firms’ inventory management. I leave these directions for future research.
3737todo: 37mention paper with Xin? Mention also papers with Wayne (ID as $q\rightarrow\infty$) and Frank/Mansi (nowcasting)? Mention them also/only in main text?
@@ A.2 NVAR($p,1$): Stationarity
3838todo: 38note: just taking largest $\lambda$ insufficient; numerically, largest eigval (in abs.v.) of AR(p) is not obtained for largest $\lambda$
@@ A.4 NVAR($p,q$)
In the previous example, the observed monthly series depends on its value in the past $p$ quarters, i.e. on its value three months ago, six months ago, etc., up to $3p$ months ago. [^48] Note that this assumes that transmission happens instantaneously at the end of each (network interaction-) period. Alternative definitions of a smoother transmission inevitably lead towards abandoning the paradigm of discrete time.
Consider next the general case: $q\in\mathbb{Q}_{++}$. We can write $q=q_{1}q_{2}$ with $q_{1}^{-1}\in\mathbb{N}$ and $q_{2}\in\mathbb{N}$. [^49] Note that $q_{2}$ is the least common multiple of $q$ and $1$, whereas $q_{1}$ is their greatest common denominator. For example, under monthly observations, $q=4/3$ implies that network interactions occur every three weeks, $q=4/5$ that they occur every five weeks, and $q=30/4$ that they occur every four days. We can model this case as a combination of the $q\in\mathbb{N}$ and $q^{-1}\in\mathbb{N}$ cases: we observe every $q_{2}\in\mathbb{N}$ periods a snapshot of an NVAR($p^{*},1)$ process with $p^{*}=p/q_{1}\in\mathbb{N}$ and with parameters restricted as in the preceding paragraph. For example, under monthly observations and $q=4/3$, this amounts to observing every fourth period a snapshot of a weekly process that depends on its value three weeks ago, six weeks ago, etc.
If $y_{t}$ is observed for $T$ periods – $t=1:T$ –, then $\tilde{y}_{\tau}$ evolves for $T_{\tau}$ (network interaction-)periods – $\tau=1:T_{\tau}$. Thereby, $T_{\tau}$ satisfies $T=|\{(1:T_{\tau})/q\}\cap\mathbb{N}|$; the number of elements in the set $1:T_{\tau}$ that are integer-multiples of $q$ equals $T$. This yields $T_{\tau}=qT$ under $q\in\mathbb{N}$, and $T_{\tau}=T$ under $q^{-1}\in\mathbb{N}$. For other $q\in\mathbb{Q}_{++}$, we have $T_{\tau}=q_{2}T$, where $q_{2}$ is least common multiple of $q$ and 11.
@@
$\underset{g\rightarrow\infty}{\lim}q\frac{\partial y_{t}}{\partial\tilde{\varepsilon}_{\tau-g}}=\underset{g\rightarrow\infty}{\lim}\sum_{h=0}^{g}\frac{\partial\tilde{y}_{\tau+g}}{\partial\tilde{\varepsilon}_{\tau+h}}\;.$
@@
which leads to $\text{Cov}(u_{it},u_{jt})=\mathbf{1}\left\{{i=j}\right\}\frac{1}{2}\sigma_{i}^{2}+\frac{1}{4}\alpha_{1}(a_{ij}\sigma_{j}^{2}+a_{ji}\sigma_{i}^{2})+\frac{1}{4}\alpha_{1}^{2}\sum_{k=1}^{n}a_{ik}a_{jk}\sigma^{2}_{k}$. In the former case, $\text{Cov}(u_{it},u_{jt})$ is determined by common exposure to third units, while in the latter case bilateral exposure matters as well. In either case, exposure means first-order connections. [^51] To understand the different results for stock and flow variables, assume for simplicity that there are no self-links: $a_{ii}=0\;\forall\;i$. For a stock variable, $y_{it}=\tilde{y}_{i,\tau}$, which implies that $u_{it}$ only depends on $\tilde{u}_{i,\tau}$, not on $\tilde{u}_{i,\tau-1}$: we have $u_{it}=\tilde{u}_{i,\tau}+\alpha_{1}\sum_{j=1}^{n}a_{ij}\tilde{u}_{j,\tau-1}$. As a result, the only (possibly) common terms in $u_{it}$ and $u_{jt}$ are the one-period lagged high-frequency innovations of third units, $\tilde{u}_{k,\tau-1}$, and any comovement between $u_{it}$ and $u_{jt}$ is due to common exposure to these units: $\text{Cov}\left({u_{it},u_{jt}}\right)\neq 0$ iff $\exists\;k$ s.t. $a_{ik},a_{jk}\neq 0$. In contrast, for a flow variable, $y_{it}=\tilde{y}_{i,\tau}+\tilde{y}_{i,\tau-1}$, which implies that $u_{it}$ also depends on $\tilde{u}_{i,\tau-1}$: we have $u_{it}=\tilde{u}_{i,\tau}+\tilde{u}_{i,\tau-1}+\alpha_{1}\sum_{j=1}^{n}a_{ij}\tilde{u}_{j,\tau-1}$. Besides common exposure to third units, comovement is also due to bilateral exposure: $a_{ij}\neq 0\;|\;a_{ji}\neq 0\;\Rightarrow\;\text{Cov}\left({u_{it},u_{jt}}\right)\neq 0$. Under $q=3$, exposure is determined by first- and second-order connections: e.g. for a stock variable, we have $u_{t}=\tilde{u}_{\tau}+\alpha_{1}A\tilde{u}_{\tau-1}+(\alpha_{2}A+\alpha_{1}^{2}A^{2})\tilde{u}_{\tau-2}$, with
@@ Appendix B NVAR: Inference
Relative to the notation used for the Kalman filter, this is with a slight abuse of notation, as $s_{\tau|\tau+1}\neq\mathbb{E}\left[{s_{\tau}|X_{1:\tau},s_{\tau+1}}\right]$ but $s_{\tau|\tau+1}=\mathbb{E}\left[{s_{\tau}|X_{1:\tau},\left[{s_{\tau+1}}\right]_{1}}\right]$, and similarly $P_{\tau|\tau+1}=\mathbb{V}\left[{s_{\tau}|X_{1:\tau},\left[{s_{\tau+1}}\right]_{1}}\right]$. See Nelson and Kim (1999, p. 194) for the adjustments of the CKSS required when $s_{t}$ is a companion form-VAR(1).
@@ C.1 Theory
$\hskip 103.50105pt\text{s.t.}\quad\quad\check{x}_{t,t-2}^{ij}$ $=x^{ij}_{t,t-2}\;,$
@@ C.3 Estimation
The Uniform prior for $\alpha\in[0,1]^{p-1}\cap\{\alpha:\sum_{l=1}^{p-1}\alpha_{l}\leq 1\}$ can be broken up as follows:3939todo: 39derivation: P009diary3,p.39
@@
For the parameters $(\sigma^{2}_{a},\sigma^{2\prime})^{\prime}\in\mathbb{R}_{++}^{n+1}$, an upper bound has to be chosen so as to specify a proper Uniform prior distribution, since draws from it are needed to initialize the SMC sampler. A high upper bound is desirable so as to avoid domain restrictions in areas with non-trivial likelihood values. However, for efficient sampling, lower values are preferred. I choose a low upper bound to initialize the sampler, but allow the algorithm to disrespect it in search for parameter-values associated with high likelihood values and therefore high posterior mass. Ex-post, I redefine the upper bound as the largest posterior draw among $(\sigma^{2}_{a},\sigma^{2\prime})^{\prime}$ across all models. The resulting MDDs could be re-scaled to take this into account, but this is not necessary, as all models’ MDD is “off” by the same proportionality constant, which means that the ranking is unaffected. [^57] Let $s=(\sigma^{2}_{a},\sigma^{2\prime})^{\prime}\in[0,\bar{s}]^{n+1}$. We get the MDD $p(Y)=\int p(Y|\theta_{-s},s)p(\theta_{-s})p(s)d(\theta_{-s},s)=\bar{s}^{-(n+1)}\int p(Y|\theta_{-s},s)p(\theta_{-s})d(\theta_{-s},s)\;.$ If the upper bound is ex-post rescaled to $s^{*}$, the MDD needs to be corrected by multiplying by $(\bar{s}/s^{*})^{n+1}$. The analogous holds with heterogeneous upper bounds. I set $\bar{\sigma}^{2}_{i}=0.5\mathbb{V}[\Delta\tilde{y}_{it}]$ and $\bar{\sigma}^{2}_{a}=\max_{i}\bar{\sigma}^{2}_{i}$.